OpenStudy (anonymous):

Use the definition of derivative as a limit to compute f'(2) for the function f(x)=square root of x+2.

OpenStudy (anonymous):

\[f'(x) = \lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

is the answer 1/4?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

\[f'(2) = \lim_{h \rightarrow 0}\frac{f(2+h)-f(2)}{h}=\lim_{h \rightarrow 0}\frac{\sqrt{4+h}-2}{h}\]